OpenStudy (anonymous):

Ichiro Suzukis batting average in 2001 was 0.350. Write and solve an equation to find George Sisters bating average in the year 1922, given that his average X was 0.070 more than Ichiro Suzuki's average in 2001.

OpenStudy (anonymous):

Can somebody please help me with this question because i'm really confused on how to solve it

OpenStudy (anonymous):

is it telling you the .070 is more to ichiro's .35 so you add togather so his average a.k.a x=.42 i am not 100% sure i have terrible english

OpenStudy (anonymous):

So 0.070 is postive so im going to subtract it wit 0.070 and then subtract -0.070 + 0.350 which gives me 0.28

OpenStudy (anonymous):

But the answers its giving me are A x-0.0350=0.070;280 B x+0.070=0.350;0.420 C x+0.070=0.350=0.280 D x-0.070=0.350;0.420

OpenStudy (anonymous):

i think it is d you are suppose get higher than ichiro's score as well .070 more which equals to .420 so i believe it is d

OpenStudy (anonymous):

Yeahh! thanks a lot Ko-Ki i appreciate it :)